I am completely confused, can you help me figure out these two:
170=[52.00+0+21.99(x-2)] +3% [52.00+0+21.99 (x-2)]
and
350=[89.99+0+19.99(x-2)] + 6% [89.99+0+19.99(x-2)]

Answer by rothauserc(4718)

1) 170 = [52.00+0+21.99(x-2)] +3% [52.00+0+21.99 (x-2)]
170 = (52.00+21.99x-43.98) + 0.03(52.00+21.99x-43.98)
170 = 21.99x+8.02 + 0.03(21.99x+8.02)
170 = 21.99x+8.02 + 0.6597x+0.2406
170 = 22.6497x + 8.2606
161.7394 = 22.6497x
x = 7.140906944
*********************
check answer by substituting for x
170 = 165.048543699 + 0.03(165.048543699)
170 = 165.048543699 + 4.951456311
170 = 170.00000001
answer checks
**********************
2) 350 = [89.99+0+19.99(x-2)] + 6% [89.99+0+19.99(x-2)]
350 = (89.99+19.99x-39.98) + 0.06(89.99+19.99x-39.98)
350 = 19.99x+50.01 + 0.06(19.99x+50.01)
350 = 19.99x+50.01 + 1.1994x + 3.0006
350 = 21.1894x + 53.0106
296.9894 = 21.1894x
x = 14.015941933
*********************
check answer by substituting for x
350 = 330.188679241 + 0.06(330.188679241)
350 = 330.188679241 + 19.811320754
350 = 349.999999995
answer checks
**********************